Decide whether or not the equation 23-x2-3x+1=0 has a root between 1 and 2. Justify your answer.


1
Expert's answer
2020-10-17T10:53:50-0400

We have function f(x)=23-x2-3x+1=-x2-3x+9.

f(1)=-1-3+9=5>0,

f(2)=-4-6+9=-1<0.

It means that it should be a root between 1 and 2.